Some problems developed today with our recent update that forced us to switch back to our primary Flash server. During the update we were running new software on our backup Flash server, but it needs more attention to work correctly.
We do everything we possibly can to avoid service interruptions during normal hours, but during this operation the streaming service was down for about 90 minutes. Service was restored at 1:45PM Pacific and is now operating normally.
If you were connected earlier today, you may need to empty your browser "cache" to reload the site (helpful instructions under EMPTY BROWSER CACHE in our HELP section). We sincerely apologize for this problem and are taking steps to prevent any recurrence.